3. (II) A charge Q is transferred from an initiall
Answer #1

F = k*Q1*Q2/r^2

Q1 = Q , Q2 = Q ,

r = 24 m = 0.24m

So, F = 9*10^9*Q^2/0.24^2 = 17*10^-3

So, Q = 3.3*10^-7 C

So, the number of elecrtons transferred = Q/e = 3.3*10^-7/(1.6*10^-19)

= 2.06*10^12